In a material science lab, you connect a 0.250 Kg air track lightweight flyer to the end of an ideal string is massless and start it oscillating. The time from when the lightweight plan initial travels trough the equilibrium point to the second time it travels through that points is 2.50 sec. find the springs constant 


1
Expert's answer
2021-02-06T10:06:33-0500

"T=2\pi\sqrt{\frac{m}{k}}\to T^2=4\pi^2\frac{m}{k}\to k=4\pi^2\frac{m}{T^2}="


"=4\cdot3.14^2\cdot\frac{0.250}{2.5^2}=0.79\ (N/m)" . Answer